OpenStudy (anonymous):

Please Help me? /.\ (i'll Fan) Thanks <3 Below are the steps to solve an equation: Step 1: |x - 5| + 2 = 5 Step 2: |x - 5| = 5 - 2 Step 3: |x - 5| = 3 Which of the following is a correct next step to solve the equation?

OpenStudy (johnweldon1993):

Well going from step 1 to step 2...yes we would subtract 2 from both sides of the equation...so it looks like step 2 is the correct way to go about this

OpenStudy (anonymous):

x + 5 = -3 -x - 5 = 3 x - 5 = -3 x + 5 = 3

OpenStudy (anonymous):

i have to choose what goes next , in order to solve /.\

OpenStudy (johnweldon1993):

Ahh thank you... |x - 5| = 3 You know that inside an absolute value bar, the answer can never be negative...so x - 5 can equal - 3...because then the absolute value bar would change that to 3... so x - 5 = -3

OpenStudy (anonymous):

thank you! im glad i found you! you used to help me on my other ones ! i had to make a new account!

OpenStudy (johnweldon1993):

Aww, Well I'm glad I could help you again :)